# Uncovering faint lensed gravitational-wave signals and reprioritizing their follow-up analysis using galaxy lensing forecasts with detected counterparts

### v2

@@ 3.1 Overview of Lensing Searches
Typically, strong lensing searches for multiple images are done in two ways. The first focuses on the GW events detected with the traditional searches. The idea is then to see if the pool of candidates contains event pairs, triplets, or quadruplets with similar frequency evolution, as expected from strong lensing. The identification is typically done in multiple steps, starting from the least expensive and less precise methods to the most precise and computationally heavier ones (Hannuksela et al., 2019; Abbott et al., 2021e, 2023; Janquart et al., 2023c). For the first filtering, two methods are used: posterior overlap (Haris et al., 2018), where one tests for the compatibility between the posteriors of parameters unaffected by lensing, and LensID (Goyal et al., 2021), a machine learning pipeline comparing the sky and time-frequency maps of the events forming the pair of interest. The GW pairs not discarded by these methods are then passed to the next step, done by GOLUM (Janquart et al., 2021a, 2023b), which uses the posteriors of the first image to analyze the strain of the second one. The remaining pairs are passed to a full joint Bayesian analysis framework also including selection and population effects to obtain a Bayes factor (Lo & Magana Hernandez, 2021) [^5] Other codes implementing joint parameter estimation exist but they generally do not include the full population and selection effect procedure (Liu et al., 2021; Janquart et al., 2023b). In this work, when doing joint parameter estimation, we use the GOLUM package (Janquart et al., 2022).. However, if some of the images are not detected by the traditional searches, these events will be missed. To remedy this problem, sub-threshold searches exist. These target GW images that are counterparts of the detected GW events by constructing a reduced template bank (Li et al., 2019; McIsaac et al., 2020; Dai et al., 2020; Li et al., 2023). For example, searching for a sub-threshold candidate for a given high-mass binary black hole could be performed by searching the GW data with only waveforms with identical frequency evolution, instead of the entire bank of all possible waveforms. This reduces the bank’s trial factor and enables one to recover fainter triggers compared to the full bank. However, such searches can lead to a relatively long list of triggers as they also pick up unrelated fainter signals or noise artefacts.
@@ 3.2 Improving the detection confidence of candidates using lensing
where $\mathcal{H}^{\text{signal}}$ and $\mathcal{H}^{\text{noise}}$ are the hypotheses that the data contain a signal or not, respectively. The first term on the right-hand side is the likelihood ratio, the second term is the prior odds, and $\mathcal{B}^{\text{signal}}_{\text{noise}}$ is the Bayes factor.
@@
It is worth noting that care would need to be taken in implementing such arrival time predictions and additional lensing information in the sub-threshold searches in practice, which is beyond the scope of this work. The Bayes factor that is often cited in the context of lensing joint parameter estimation is derived after applying extensive Bayesian analyses methods. For searches, the information is found more rapidly by looking through a template bank, without using the signal-versus-noise Bayes factor to quantify the detection confidence. Instead, one uses different statistics to describe the likelihood of the trigger being a GW signal against it being noise. This is based on a likelihood ratio for the GstLAL pipeline (Li et al., 2019, 2023), or other ranking statistics for the PyCBC ones (McIsaac et al., 2020; Nitz et al., 2020; Davies et al., 2020). This statistic is the same between dedicated lensing searches and usual unlensed ones. We also note the possibility of having a difference in data quality between the data used for searches and parameter estimation analyses since more in-depth data cleaning is applied to the latter if needed. Adding lensing-based information will enhance the confidence in the lensing status of detected triggers. For these reasons, this work investigates how much the arrival time will aid in identifying and re-ranking triggers, to understand how important an implementation of accurate arrival time predictions is. The realistic implementation of this method in the searches is beyond the scope of this work. The next section explains how we can build such a distribution and how it can be used to re-rank the sub-threshold triggers.
@@ 4.1 The Lensed Time Delay Distribution
Suppose one detects two or three super-threshold lensed images, with one or more sub-threshold images that have been missed by standard searches, to predict the arrival time of the sub-threshold images. In that case, one can analyze the super-threshold events using JPE (Liu et al., 2021; Lo & Magana Hernandez, 2021; Janquart et al., 2021a, 2023b). The procedure gives one access to (i) the Morse factor difference, generally seen as a discrete phase difference between two images, (ii) the difference in arrival times, obtained with millisecond precision, and (iii) the relative magnifications, less precisely determined as their error are correlated with the luminosity distance uncertainty. Since the relative magnifications are subject to more uncertainty, they tend to be less relevant in predicting the arrival times of the sub-threshold events. Nevertheless, for completeness, we use all three relative lensing parameters to predict the time delay distributions.
Using the Morse factor, arrival time, and relative magnification differences, the goal is to find at what time any sub-threshold images arrive in the data. That is, to forecast the arrival time of the sub-threshold event using the time-delay prior:
@@
where $\vec{d}_{\text{super}}$ are all the gravitational-wave data from the super-threshold events, $t_{s1}$ is the time delay between the sub-threshold trigger and the super-threshold image arriving first, $\vec{\Phi}_{j1}$ [^6] From here on, the sub-scripts correspond to the ordering of the detected super-threshold images and not the position of the images in the lensed multiplet. So, for example, $\vec{\Phi}_{21}$ corresponds to the relative parameters between the first and second detected super-threshold images. are the relative lensing parameters between the first and all other (second and/or third) super-threshold images —depending on the lens-source system— and $\mathcal{M}$ represent the assumed BBH and lens models. The posterior distribution $p(\vec{\Phi}_{j1}|\vec{d}_{\text{super}},\mathcal{H}_{\mathrm{L}},\mathcal{M})$ are thus the measured lensing parameters from the super-threshold gravitational-wave data.
@@
where now $\vec{\Phi}_{j1}$ are understood to be the median. From this point onwards, we will make this approximation and drop the explicit mention of the GW data and posterior. If we select a given trigger and evaluate Eq. (7), we obtain the probability to observe a trigger with a specific time delay given the characteristics of the super-threshold events and the assumed lens and source models. The process to do so is illustrated in Fig. 2.
@@
To construct the time delay prediction (the probability density function described in Eq. (7)), we rely on a Gaussian kernel density estimator (KDE) constructed based on a large catalogue of lensed events. The latter is generated using the LeR package (Phurailatpam, 2023), an efficient code to generate lensing statistics. Principally, the code generates the catalogue by sampling the characteristics of the sources and lenses from chosen source and lens populations, then solving the lens equation and retaining only the setups leading to multiple images. Here, we use the PowerLaw+Peak BBH mass model (Abbott et al., 2021c), the redshift distribution from Oguri (2018), and an elliptical power-law with shear lens model (Tessore & Benton Metcalf, 2015). For all the models, we use the same parameters as the one specified in Wierda et al. (2021). Here, and throughout the work, we use the detector network made of the two LIGO and the Virgo detectors and assume they are at their expected O4 sensitivity (Abbott et al., 2020). When selecting the samples on which we apply the KDE, we consider as super-threshold events those with a network signal-to-noise-ratio (SNR) higher than 8 and sub-threshold those with a network SNR between 6 and 8. The value of 8 is chosen because, in Gaussian stationary noise, it is a good proxy for detectability (Essick, 2023). For the sub-threshold case, the SNR values are more uncertain but we take an SNR cut at 6 to remove the faintest signals, which would be very hard to detect even with adapted search methods. Using LeR, we generate a catalogue containing $2\times 10^{6}$ lensed systems. This number is chosen to ensure we have enough samples representing the possible lensing scenarios while keeping a reasonable generation time for the catalogue.
@@
[…]
@@
In the above example, it is worth noting that since the Morse factor differences are sufficient to tell we have seen the two first images, we also know that no genuine counterpart can exist before the second observed image, meaning the time delay probability is zero and we can exclude any trigger falling in that region. This effect enables us to reduce the list of triggers to follow-up. A region with a zero probability always arises when we have three images. When we have only two images, such a region can exist only if we have two images with the same type, as we know no image should exist in between. However, in this case, the region is generally shorter and the list reduction is milder. This can also be observed in Fig. 3, where we represent the final distribution obtained depending on the observed super-threshold images and covering the cases depicted in Table 1. A more detailed breakdown with the various contributions of each sub-case is given in Appendix A. In particular, Figs 3b3e, and 3f show the cases where we have a zero probability region.
@@
Besides, Fig. 3a shows the time delay distribution when a single super-threshold image is observed. The probability of the sub-threshold image arriving later is larger than before. It is expected since the third and fourth images are more likely to be sub-threshold. This result is similar to that presented in other work (Wierda et al., 2021; More & More, 2022; Goyal et al., 2023). For all the figures, discontinuities at the super-threshold events can be observed. This behaviour is expected as observing a sub-threshold event before and after a super-threshold one corresponds to different situations with different probabilities. Figs. 3c and  3d, also show that when one has two super-threshold images with different image types, i.e. $n_{21}=0.5$, we do not have a region with a clear zero probability, meaning no triggers will be immediately discarded. Still, the features of the probability distribution positively impact the re-ranking process.
@@ 4.2 Re-Ranking the Triggers
Above, we showed the behaviour of the time delay distribution under the lensed hypothesis. Here, we go through how it can be used to re-rank triggers. One can use an approach similar to the ones used in Haris et al. (2018); Wierda et al. (2021); More & More (2022); Janquart et al. (2023a); Goyal et al. (2023), where the probability under the lensed hypothesis is compared to that under the unlensed hypothesis in the ratio
@@
Once we also have the unlensed KDE, we can compute $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ for all the triggers corresponding to a given set of lensed super-threshold images by evaluating the lensed and unlensed distributions. In the cases where the lensed probability is zero, $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}=0$, and triggers are directly discarded. For the others, we can rank them in descending order of $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$, obtaining a ranked list of triggers. The time delay of a lower-ranked trigger is less consistent with the lensed hypothesis than the ones with a higher rank. Since this method is statistical, there can be cases where the lensed counterpart is in a lower probability region than some unlensed triggers. Therefore, this does not guarantee the genuine counterpart to always end up at the top rank. However, the method suggested here is complementary to other methods such as using sky map overlap between the triggers and the super-threshold images (Wong et al., 2021; Goyal et al., 2023) —which should be more efficient when considering more than one super-threshold image thanks to the reduced sky location offered by the joint observation of multiple images (Janquart et al., 2021a)— or using the proximity in chirp mass (Goyal et al., 2023). Studying the full interplay between those methods and our approach is left for future work.
@@ 5.1 Mock Data Setup
In this section, we demonstrate our method’s ranking capabilities on an approximate experimental setup. We generate 50 lensed systems with two or three super-threshold events and at least one sub-threshold event, following the population described above. Our testing set has 30 cases with two super-threshold images and 20 where we could observe a super-threshold triplet. We perform JPE on all these doubles and triples using the GOLUM package (Janquart et al., 2022) and use their posteriors to evaluate $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$. More precisely, we use the median value for the time delay and the relative magnification, and the dominant mode of the Morse factor difference.
To avoid the computational burden of simulating a full year of data in which we inject each image of the lensed systems, we proceed more approximately. We generated 20000 unlensed events distributed over the year with LeR and the same BBH population models in Sec.4 and keep only the sub-threshold ones, hence with their network SNR between 6 and 8. This leads to 132 sub-threshold false triggers for each lensed system. Since we do not run the sub-threshold searches, we cannot associate the usual likelihood ratio and FAR as the ranking statistic. Instead, we use an approach closer to representing what would be done if some rapid follow-up has already been done by ranking the events according to their proximity in chirp mass (Goyal et al., 2023). However, since we do not do template bank searches, we do not have access to the best-matching template. Therefore, we use the posterior overlap approach, computing the associated Bayes factor (Haris et al., 2018) using only the chirp mass posteriors, noted $\mathcal{B}_{\mathcal{M}_{c}}$. Moreover, to avoid the computational burden of running parameter estimation on all the sub-threshold triggers, including the lensed counterparts and the unlensed events, we generate a mock posterior by assuming it is a Gaussian distribution with a mean equal to the true value of the chirp mass, and standard deviation equal to the mean mass divided by the SNR, following the example from Çalışkan et al. (2023). For the super-threshold images, we use the posteriors coming from JPE. We then use the $\mathcal{B}_{\mathcal{M}_{c}}$ statistics for the various events as an initial ranking.

### v3

@@ 3.1 Overview of Lensing Searches
Typically, strong lensing searches for multiple images are done in two ways. The first focuses on the GW events detected with the unlensed searches. The idea is then to see if the pool of candidates contains event multiplets with similar frequency evolution, as expected from strong lensing. The identification is typically done in multiple steps, starting from the least expensive and less precise methods to the most precise and computationally heavier ones (Hannuksela et al., 2019; Abbott et al., 2021e, 2023; Janquart et al., 2023c). For the first filtering, two methods are used: posterior overlap (Haris et al., 2018), where one tests for the compatibility between the posteriors of parameters unaffected by lensing, and LensID (Goyal et al., 2021), a machine learning pipeline comparing the sky and time-frequency maps of the events forming the pair of interest. The GW pairs not discarded by these methods are then passed to the next step, done by GOLUM (Janquart et al., 2021a, 2023b), which uses the posteriors of the first image to analyze the strain of the second one. The remaining pairs are passed to a full joint Bayesian analysis framework also including selection and population effects to obtain a Bayes factor (Lo & Magana Hernandez, 2021) [^5] Other codes implementing joint parameter estimation exist but they generally do not include the full population and selection effect procedure (Liu et al., 2021; Janquart et al., 2023b). In this work, when doing joint parameter estimation, we use the GOLUM package (Janquart et al., 2022).. However, if some of the images are not detected by the traditional searches—typically because they are too faint, these events will be missed. Sub-threshold searches exist to remedy this problem. These target GW images that are counterparts of the detected GW events by constructing a reduced template bank (Li et al., 2019; McIsaac et al., 2020; Dai et al., 2020; Li et al., 2023). This reduces the bank’s trial factor and enables one to recover fainter triggers compared to the full bank. However, such searches can lead to a relatively long list of triggers as they also pick up unrelated fainter signals or noise artefacts.
@@ 3.2 Improving the detection confidence of candidates using lensing
where $\mathcal{H}^{\text{signal}}$ and $\mathcal{H}^{\text{noise}}$ are the hypotheses that the data contain a signal or not, respectively. The first term on the right-hand side is the likelihood ratio, also called the Bayes factor ($\mathcal{B}^{\text{signal}}_{\text{noise}}$), and the second term is the prior odds.
@@
It is worth noting that care would need to be taken in implementing such arrival time predictions and additional lensing information in the sub-threshold searches in practice, which is beyond the scope of this work. The Bayes factor that is often cited in the context of lensing joint parameter estimation is derived after applying extensive Bayesian analyses methods. For searches, the information is found by looking through a template bank. Instead, one uses different statistics to describe the likelihood of the trigger being a GW signal against it being noise. This is based on a likelihood ratio for the GstLAL pipeline (Li et al., 2019, 2023), or other ranking statistics for the PyCBC ones (McIsaac et al., 2020; Nitz et al., 2020; Davies et al., 2020). This statistic is the same between dedicated lensing searches and usual unlensed ones. We also note the possibility of having a difference in data quality between the data used for searches and parameter estimation analyses since more in-depth data cleaning is applied to the latter if needed. Adding lensing-based information will enhance the confidence in the lensing status of detected triggers. For these reasons, this work investigates how much the arrival time will aid in identifying and re-ranking triggers to understand importance of an implementation of accurate arrival time predictions. The realistic implementation of this method in the searches is beyond the scope of this work.
@@ 4.1 The Lensed Time Delay Distribution
Suppose one detects two or three super-threshold lensed images, with one or more sub-threshold images that have been missed by standard searches, to predict the arrival time of the sub-threshold images. In that case, one can analyze the super-threshold events using JPE (Liu et al., 2021; Lo & Magana Hernandez, 2021; Janquart et al., 2021a, 2023b). The procedure gives one access to (i) the Morse factor difference, generally seen as a discrete phase difference between two images, (ii) the difference in arrival times, obtained with millisecond precision, and (iii) the relative magnifications, less precisely determined as their error are correlated with the luminosity distance uncertainty. Since the relative magnifications are subject to higher uncertainty, they tend to be less relevant in predicting the arrival times of the sub-threshold events. Nevertheless, for completeness, we use all three relative lensing parameters to predict the time delay distributions.
Using the Morse factor, arrival time, and relative magnification differences, the goal is to find the arrival time of any sub-threshold images in the data. That is, to forecast the arrival time of the sub-threshold event using the time-delay prior:
@@
where $\vec{d}_{\text{super}}$ are all the gravitational-wave data from the super-threshold events, $t_{s1}$ is the time delay between the sub-threshold trigger and the super-threshold image arriving first, $\vec{\Phi}_{j1}$ [^6] From here on, the sub-scripts correspond to the ordering of the detected super-threshold images and not the position of the images in the lensed multiplet. So, for example, $\vec{\Phi}_{21}$ corresponds to the relative parameters between the first and second detected super-threshold images. are the relative lensing parameters between the first and all other (second and/or third) super-threshold images—depending on the lens-source system—and $\mathcal{M}$ represent the assumed BBH and lens models, which can be chosen when doing the analysis. The posterior distribution $p(\vec{\Phi}_{j1}|\vec{d}_{\text{super}},\mathcal{H}_{\mathrm{L}},\mathcal{M})$ are thus the measured lensing parameters from the super-threshold GW data.
@@
where now $\vec{\Phi}_{j1}$ are understood to be the median. From this point onwards, we will make this approximation and drop the explicit mention of the GW data and posterior, though the method can be applied to all points in the posteriors to obtained a final distribution as output. If we select a given trigger and evaluate Eq. (7), we obtain the probability to observe a trigger with a specific time delay given the characteristics of the super-threshold events and the assumed lens and source models. The process is illustrated in Fig. 2.
@@
To construct the time delay prediction (the probability density function described in Eq. (7)), we rely on a Gaussian kernel density estimator (KDE) constructed based on a large catalogue of lensed events. The latter is generated using the LeR package (Phurailatpam, 2023), an efficient code to generate lensing statistics. Principally, the code generates the catalogue by sampling the characteristics of the sources and lenses from chosen source and lens populations, then solving the lens equation and retaining only the setups leading to multiple images. Here, we use the PowerLaw+Peak BBH mass model (Abbott et al., 2021c), the redshift distribution from Oguri (2018), and an elliptical power-law with shear lens model (Tessore & Benton Metcalf, 2015). For all the models, we use the same parameters as the one specified in Wierda et al. (2021). The parameters are also given in Appendix B. Here, and throughout the work, we use the detector network made of the two LIGO and the Virgo detectors and assume they are at their expected O4 sensitivity (Abbott et al., 2020). When selecting the samples on which we apply the KDE, we consider as those with a network signal-to-noise-ratio (SNR) higher than 8 super-threshold events and those with a network SNR between 6 and 8 as sub-threshold. The value of 8 is chosen because, in Gaussian stationary noise, it is a good proxy for detectability (Essick, 2023). For the sub-threshold case, the SNR values corresponding to identifiable signals are more uncertain, but here, we take an SNR cut at 6 to remove the faintest signals, which would be very hard to detect even with adapted search methods. Using LeR, we generate a catalogue containing $2\times 10^{6}$ lensed systems. This number is chosen to ensure having enough samples representing the possible lensing scenarios while keeping a reasonable generation time for the catalogue.
@@
Let us illustrate the process with a simple example where we have observed three super-threshold images with $n_{21}=0$ and $n_{31}=0.5$ (4th row in Table 1). This situation has two sub-cases: one where the missing image is the fourth one in the quadruplet and one where it is the third image. So, we make the first KDE by selecting events in the catalogue with the three first images super-threshold, and the last one sub-threshold, according to the SNR criterion defined earlier. Then, we do the same for the case where the third image is sub-threshold and all the others have their SNR larger than 8. With the catalogue, we obtained 389 and 52 lensed events matching the first and second cases respectively. [^7] The difference in the sample sizes is reasonable because the fourth image tends to be fainter than the third, leading to a larger sample size for the first case. Using these samples, we then make two KDEs representing the two probability distributions
@@
In the above example, it is worth noting that since the Morse factor differences are sufficient to tell we have seen the two first images, we also know that no genuine counterpart can exist before the second observed image, meaning the time delay probability is zero and we can exclude any trigger falling in that region. A region with a zero probability always arises when we have three images. When we have only two images, such a region can exist only if we have two images with the same type, as we know no image should exist in between. However, in this case, the region is generally shorter and the reduction in triggers to follow-up is milder. This can also be observed in Fig. 3, where we represent the final distribution obtained depending on the observed super-threshold images and covering the cases depicted in Table 1. A more detailed breakdown with the various contributions of each sub-case is given in Appendix A. In particular, Figs 3b3e, and 3f show the cases where we have a zero probability region.
@@
For reference, Fig. 3a shows the time delay distribution when a single super-threshold image is observed. The probability of the sub-threshold image arriving later is larger than before. It is expected since the third and fourth images are more likely to be sub-threshold. This result is similar to that presented in other work (Wierda et al., 2021; More & More, 2022; Goyal et al., 2023). For all the figures, discontinuities at the super-threshold events can be observed. This behaviour is expected as observing a sub-threshold event before and after a super-threshold one corresponds to different situations with different probabilities. Figs. 3c and  3d, also show that when one has two super-threshold images with different image types, i.e. $n_{21}=0.5$, we do not have a region with a clear zero probability, meaning no triggers will be immediately discarded. Still, the features of the probability distribution positively impact the re-ranking process.
@@ 4.2 Re-Ranking the Triggers
Above, we showed the behaviour of the time delay distribution under the lensed hypothesis. Here, we describe how it can be used to re-rank triggers. One can use an approach similar to the ones used in Haris et al. (2018); Wierda et al. (2021); More & More (2022); Janquart et al. (2023a); Goyal et al. (2023), where the probability under the lensed hypothesis is compared to that under the unlensed hypothesis in the ratio
@@
Once we also have the unlensed KDE, we can compute $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$ for all the triggers corresponding to a given set of lensed super-threshold images. In the cases where the lensed probability is zero, $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}=0$, and triggers are directly discarded. For the others, we can rank them in descending order of $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$, obtaining a ranked list of triggers. The time delay of a lower-ranked trigger is less consistent with the lensed hypothesis than the ones with a higher rank. Since this method is statistical, there can be cases where the lensed counterpart is in a lower probability region than some unlensed triggers. However, the method suggested here is complementary to other methods such as using sky map overlap between the triggers and the super-threshold images (Wong et al., 2021; Goyal et al., 2023)—which should be more efficient when considering more than one super-threshold image thanks to the reduced sky location offered by the joint observation of multiple images (Janquart et al., 2021a)—or using the proximity in chirp mass (Goyal et al., 2023). Studying the full interplay between those methods and our approach is left for future work.
@@ 5.1 Mock Data Setup
In this section, we demonstrate our method’s ranking capabilities on an approximate experimental setup. We generate 50 lensed systems with two or three super-threshold events and at least one sub-threshold event, following the population described above. Our testing set has 30 cases with two super-threshold images and 20 where we could observe a super-threshold triplet. We perform JPE on all these doubles and triples using the GOLUM package (Janquart et al., 2022) and use their posteriors to evaluate $\mathcal{R}_{\mathrm{gal}}^{\mathrm{sub}}$. More precisely, we use the median values for the time delay and the relative magnification, and the dominant modes of the Morse factor difference. The Morse factor differences is expected to be recovered especially confidently for genuine lensed events due to its discrete nature.
To avoid the computational burden of simulating a full year of data in which we inject each image of the lensed systems, we proceed more approximately. We generated 20000 unlensed events distributed over the year with LeR and the same BBH population models in Sec.4 and keep only the sub-threshold ones, hence with their network SNR between 6 and 8. This leads to 132 sub-threshold false triggers for each lensed system. Since we do not run the sub-threshold searches, we cannot associate the usual likelihood ratio and FAR as the ranking statistic. Instead, we use an approach closer to representing what would be done if some rapid follow-up has already been done by ranking the events according to their proximity in chirp mass (Goyal et al., 2023). However, since we do not do template bank searches, we do not have access to the best-matching template. Therefore, we use the posterior overlap approach, computing the associated Bayes factor (Haris et al., 2018) using only the chirp mass posteriors, noted $\mathcal{B}_{\mathcal{M}_{c}}$. Moreover, to avoid the computational burden of running parameter estimation on all the sub-threshold triggers, including the lensed counterparts and the unlensed events, we generate a mock posterior by assuming it is a Gaussian distribution with a mean equal to the true value of the chirp mass, and standard deviation equal to the mean mass divided by the SNR, following Çalışkan et al. (2023). For the super-threshold images, we use the posteriors coming from JPE. We then use the $\mathcal{B}_{\mathcal{M}_{c}}$ statistics for the various events as an initial ranking.